Exercise 3: CHANGE IN ESTIMATE (HOMEWORK) ABC Corp. purchased equipment for $1,020,000 which was estimated to have a useful life of 16 years with a salvage value of $20,000 at the end of that time. Depreciation has been recorded for 10 years on a straight-line basis. In 2018 (year 11), it is determined that the total estimated life should be 20 years with a salvage value of $25,000 at the end of that time. Compute depreciation amount for the year 2018.
Answer #1

Depreciation per year = (Cost - Salvage value)/Useful life

= (1,020,000 - 20,000)/16 = 62,500 per year

Accumulated Depreciation upto 2018

= 62,500 * 10 years = 625,000

Book value at the beginning of 2018 = 1,020,000 - 625,000

= 395,000

Now, estimated life is changed to 20 years (only 10 years remaining as 10 years have already passed) and salvage value is changed to 25,000

Depreciation for 2018 = (395,000 - 25,000)/10

= 37,000